absolute value of x-1 -0.5 is less than or equal to (1/2)x+5
graph with points

Answer by stanbon(75887)

absolute value of x-1 -0.5 is less than or equal to (1/2)x+5
--------------
|x-1|-0.5 <=(1/2)x+5
|x-1|<=(1/2)x+5.5
--------------------
Comment: The absolute can have two meanings:
If x>=1 |x-1|=x-1
If x<1 |x-1|=-(x-1)
---------------------
So,
If x>=1, you get x-1<=(1/2)x+5.5
(1/2)x<=6.5
x<=13
-----------
If x<1, you get -(x-1)<=(1/2)x+5.5
-x+1<=(1/2)x+5.5
(3/2)x>=-4.5
x>=-3
--------------
Final solution:
-3<=x<=13
Cheers,
Stan H.